OpenStudy (anonymous):

Use a graph and a table to solve the system. Check your answer. Enter your answer as an ordered pair, like this: (3, -12) 3x+y=-12 x+3y=4

OpenStudy (anonymous):

Make a table like so: \[\begin{array}{|c|c|} \hline x &y\\ \hline 1&?\\ \hline 0&?\\ \hline -1&?\\ \hline \end{array}\] \[3x+y=-12 \Rightarrow y = -12 - 3x\] Sub in the \(x\) values. Record the \(y\) values, then plot the points. Draw a line through these points, the line should be linear. Do the same for the second equation, where the lines intersect is the solution.

OpenStudy (anonymous):

@Opcode² how?

OpenStudy (anonymous):

You see how I converted \(3x + y = -12\) to \(y = mx + d\) form? \[3x + y = -12 \Rightarrow y = -3x - 12\] In the table I gave I listed random values for \(x\). Simple sub in those values to get the value of \(y\) that corresponds with \(x\). Say: \[x = 1\] \[y = -3(1) - 12\] \[y = -3 - 12\] \[y = -15\] So when \(x\) equals 1, \(y\) equals -15. Do this a few more times till you can see the linear pattern in the points then draw a line through them, do this for the second equation as well. Were the lines intersect (cross) that is your solution.)